OpenStudy (anonymous):

5-4/5x=-11? would my answer be x=20

OpenStudy (anonymous):

please use brackets!

OpenStudy (anonymous):

\[\huge 5-\frac{4}{5}X=-11\] \[\huge \frac{-4}{5}X=-11-5\] \[\huge \frac{-4}{5}X=-16\] \[\huge X=-16 * \frac{-5}{4} =20\]